# Omnimodal Understanding for Multimodal LLMs

### v2

@@ Abstract
[…]
@@ 1 Introduction
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [51, 25, 50, 59, 61, 79, 44, 12] have led to remarkable progress on a series of tasks, for example, classification [37], captioning [1, 10, 36], question-answering [58, 52, 41], OCR [47, 80], segmentation [26, 69, 23], autonomous driving [48, 54, 4] and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have started incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench [32] and AV-Odyssey Bench [20] mainly emphasize image evaluation, whereas other benchmarks [19, 30, 74] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
@@
We conduct extensive evaluations for a broad spectrum of MLLMs, including open-source video models, video-audio models, and proprietary systems. Results reveal significant limitations in current models’ ability to reason over omni-modal inputs in real-world contexts. Specifically, open-source video-audio models, despite processing both modalities, achieve only 25% accuracy—comparable to random guessing. In contrast, proprietary models such as Gemini 1.5 Pro reach up to 48% accuracy. However, when restricted to a single modality (audio or video), Gemini’s performance drops by 15%, highlighting the critical role of integrated modality processing.
@@
To summarize, we have made the following contributions: (i) we present WorldSense, the first benchmark tailored for evaluating MLLMs’ ability on omni-modal video understanding, characterized by integrated audio-visual inputs, diverse content, and high-quality question-answering annotations; (ii) we have conducted extensive evaluation of existing MLLMs, showing that most open-source models perform near chance, and even the best proprietary model achieves only 48% accuracy—exposing a significant gap in real-world omni-modal reasoning; (iii) through ablation and failure analysis, we identify the key factors influencing performance, including raw audio and visual cues, and provide actionable insights to guide future omni-modal understanding design.
@@ 2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as [79, 36, 84, 11, 66, 53, 45, 12], successfully combine vision and text modalities. Subsequent research extends to temporal understanding [68, 25, 60, 40, 63, 67, 27, 14, 72, 78, 62, 6, 42, 35, 73], while parallel efforts [57, 9, 8] focus on audio processing. Recently, researchers shift attention to models [7, 56, 60, 43, 61] capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
Multimodal Benchmarks. The development of MLLMs has been driven by benchmarks, evolving from static image understanding [81, 38, 29, 28, 16, 77] to temporal comprehension [31, 39, 55, 83, 13, 17, 24, 65, 70, 76, 34, 3]. However, these benchmarks largely overlook the crucial role of audio in real-world perception. While several audio-visual benchmarks have been proposed, they face significant limitations. AV-Odyssey Bench [20] and OmniBench [32] focus on static images, Music-AVQA [30] and AVQA [74] are domain-specific with monotonous questions, and LongVALE [19] limits its assessment to captioning capabilities alone. Given that existing benchmarks fail to provide a comprehensive evaluation of MLLMs’ real-world understanding capabilities, we introduce WorldSense to address this critical gap in the field.
@@ 3.2 Data Collection & Curation
We primarily source our video content from FineVideo [15], a large-scale dataset comprising high-quality YouTube videos that exhibit strong audio-visual correlations across diverse real-world scenarios. To enrich the benchmark’s coverage of musical content, we supplement it with selected videos from MusicAVQA [30], ensuring a more balanced representation of auditory modalities.
@@ 3.3 Annotation Protocol
We also use MLLMs for automated verification. Vision-language models like Qwen2-VL[63] verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2[7] and OneLLM [22] are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
@@ 3.4 Dataset Statistics
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi task Open domain Sub./ Aud. A-V Correlations MSRVTT-QA [70] V 2,990 15.2 72,821 A 8.4 ActivityNet-QA [76] V 800 111.4 8,000 M 10.2 MVBench [31] V 3,641 16.0 4,000 A 27.3 MovieChat [55] V 130 500.0 1,950 M - Video-Bench [49] V 5,917 56.0 17,036 A&M 21.3 EgoSchema [46] V 5,063 180.0 5,063 A&M 126.8 Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 MMBench-Video [13] V 609 165.4 1,998 M 19.3 AVQA [74] A+V 57,000 10 57,335 M 14.2 Music-AVQA [30] A+V 9,288 60 45,867 M 8.6 OmniBench [32] A+I 1,142 M 37.8 AV-Odyssey [20] A+I 4,555 M 19.5 LongVALE [19] A+V 8,400 235 A&M WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2
WorldSense features diverse audio types such as speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [20], OmniBench [32]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first to comprehensively evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a comprehensive benchmark for real-world multimodal understanding assessment.
@@ 4.1 Settings
To comprehensively assess the multi-modal understanding ability, we evaluate three types of MLLMs: (1) open-source audio-visual models, such as Unified-IO-2 [43], OneLLM [22], and VideoLLaMA2 [7]; (ii) open-source MLLMs, such as Qwen2-VL [64], LLaVA-OneVision [27], InternVL2.5 [5], LLaVA-Video [82], and so on; (iii) proprietary MLLMs, such as Claude 3.5 Sonnet [2], GPT 4o [25], and Gemini 1.5 Pro [60]. For all evaluations, we strictly adhere to each model’s official implementation guidelines and the recommended pre-processing procedures. Video frame extraction follows the official configurations specified by corresponding MLLMs, while proprietary models are evaluated according to their API specifications and recommended input formats. Model performance is assessed through direct comparison between model outputs and ground-truth answers.
@@
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [43] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [43] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [43] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [22] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 VITA-1.5 [18] 7B 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9 Qwen2.5-Omni [71] 7B 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4 Open-Source Video MLLMs Video-LLaVA [33] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [21] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [64] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [75] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [27] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [82] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [25] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [60] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
@@ 4.2 Results on WorldSense
[…]
@@
[…]
@@
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Unified-IO-2 L [43] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [43] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [43] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [22] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 VITA-1.5 [18] Audio 30.2 35.6 36.3 30.9 32.2 32.2 31.4 33.3 32.9 + Caption 39.2 39.8 37.2 37.5 37.5 35.2 34.9 38.4 37.5+4.6 + Video 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9+4.0 Qwen2.5-Omni [71] Audio 40.0 38.2 36.0 33.5 31.1 30.5 32.3 33.3 34.9 + Caption 40.0 37.9 38.9 33.5 36.7 37.8 37.7 38.9 37.9+3.0 + Video 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4+10.5 Gemini 1.5 Pro [60] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4

### v3

# arXiv:2502.04326v3[cs.CV] 01 Mar 2026

@@ Abstract
We introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs. In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality. Based on our WorldSense, we extensively evaluate various state-of-the-art models. The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ($65.1\%$ best accuracy). By analyzing the limitations of current models, we aim to provide valuable insight to guide development of real-world understanding. We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
@@ 1 Introduction
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) (OpenAI, 2023; Hurst et al., 2024; OpenAI, ; Team et al., 2023, 2024b; Zhang et al., 2023; Ma et al., 2024a; Fang et al., 2023) have led to remarkable progress on a series of tasks, for example, classification (Liu et al., 2024c), captioning (Alayrac et al., 2022; Dai et al., 2023; Liu et al., 2024b), question-answering (Tang et al., 2024; Panagopoulou et al., 2023; Liu et al., 2024f), OCR (Mathew et al., 2021; Zhang et al., 2024b), segmentation (Lai et al., 2024; Xia et al., 2024; He et al., 2024a), autonomous driving (Nie et al., 2025; Sima et al., 2025; Chen et al., 2024a) and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have started incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench (Li et al., 2024d) and AV-Odyssey Bench (Gong et al., 2024) mainly emphasize image evaluation, whereas other benchmarks (Geng et al., 2024; Li et al., 2022; Yang et al., 2022) either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
@@
We conduct extensive evaluations for a broad spectrum of MLLMs, including open-source video models, video-audio models, and proprietary systems. Results reveal significant limitations in current models’ ability to reason over omni-modal inputs in real-world contexts. Specifically, open-source video-audio models, despite processing both modalities, achieve only 25% accuracy—comparable to random guessing. In contrast, proprietary models such as Gemini 2.5 Pro reach up to 65.1% accuracy. However, when restricted to a single modality (audio or video), existing model’s performance drops greatly, highlighting the critical role of integrated modality processing.
@@
To summarize, we have made the following contributions: (i) we present WorldSense, the first benchmark tailored for evaluating MLLMs’ ability on omni-modal video understanding, characterized by integrated audio-visual inputs, diverse content, and high-quality question-answering annotations; (ii) we have conducted extensive evaluation of existing MLLMs, showing that most open-source models perform near chance, and even the best proprietary model achieves only 65% accuracy—exposing a significant gap in real-world omni-modal reasoning; (iii) through ablation and failure analysis, we identify the key factors influencing performance, including raw audio and visual cues, and provide actionable insights to guide future omni-modal understanding design.
@@ 2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as (Zhang et al., 2023; Liu et al., 2024b; Zhu et al., 2023; Driess et al., 2023; Wang et al., 2024d; Pi et al., 2023; Ma et al., 2024b; Fang et al., 2023), successfully combine vision and text modalities. Subsequent research extends to temporal understanding (Wang et al., 2024f; Hurst et al., 2024; Team et al., 2024a; Liu et al., 2024e; Wang et al., 2024b, e; Li et al., 2024a; Fang et al., 2024b; Xu et al., 2024; Zhang et al., 2024a; Tong et al., 2024; Chen et al., 2024c; Lu et al., 2024a; Liu et al., 2024a; Yan et al., 2025; Hong et al., 2025; Zheng et al., 2025), while parallel efforts (Tang et al., 2023; Chu et al., 2023, 2024) focus on audio processing. Recently, researchers shift attention to models (Cheng et al., 2024; Sun et al., 2024; Team et al., 2024a; Lu et al., 2024b; Team et al., 2024b) capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
Multimodal Benchmarks. The development of MLLMs has been driven by benchmarks, evolving from static image understanding Zhang et al. (2024c); Liu et al. (2025); Li et al. (2023, 2024b); Fu et al. (2024a); Yue et al. (2024) to temporal comprehension (Li et al., 2024c; Liu et al., 2024d; Song et al., 2024; Zhou et al., 2024; Fang et al., 2024a; Fu et al., 2024b; He et al., 2024b; Wang et al., 2024c; Xu et al., 2017; Yu et al., 2019; Lin et al., 2024; Chandrasegaran et al., 2024). However, these benchmarks largely overlook the crucial role of audio in real-world perception. While several audio-visual benchmarks have been proposed, they face significant limitations. AV-Odyssey Bench (Gong et al., 2024) and OmniBench (Li et al., 2024d) focus on static images, Music-AVQA (Li et al., 2022) and AVQA (Yang et al., 2022) are domain-specific with monotonous questions, and LongVALE (Geng et al., 2024) limits its assessment to captioning capabilities alone. Given that existing benchmarks fail to provide a comprehensive evaluation of MLLMs’ real-world understanding capabilities, we introduce WorldSense to address this critical gap in the field.
@@ 3.2 Data Collection & Curation
We primarily source our video content from FineVideo (Farré2024FineVideo), a large-scale dataset comprising high-quality YouTube videos that exhibit strong audio-visual correlations across diverse real-world scenarios. To enrich the benchmark’s coverage of musical content, we supplement it with selected videos from MusicAVQA (Li et al., 2022), ensuring a more balanced representation of auditory modalities.
@@ 3.3 Annotation Protocol
We also use MLLMs for automated verification. Vision-language models like Qwen2-VL(Wang et al., 2024b) verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2(Cheng et al., 2024) and OneLLM (Han et al., 2024) are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
@@ 3.4 Dataset Statistics
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi task Open domain Sub./ Aud. A-V Correlations MSRVTT-QA (Xu et al., 2017) V 2,990 15.2 72,821 A 8.4 ActivityNet-QA (Yu et al., 2019) V 800 111.4 8,000 M 10.2 MVBench (Li et al., 2024c) V 3,641 16.0 4,000 A 27.3 MovieChat (Song et al., 2024) V 130 500.0 1,950 M - Video-Bench (Ning et al., 2023) V 5,917 56.0 17,036 A&M 21.3 EgoSchema (Mangalam et al., 2023) V 5,063 180.0 5,063 A&M 126.8 Video-MME (Fu et al., 2024b) V 900 1017.9 2,700 M 35.7 3086.5 MMBench-Video (Fang et al., 2024a) V 609 165.4 1,998 M 19.3 AVQA (Yang et al., 2022) A+V 57,000 10 57,335 M 14.2 Music-AVQA (Li et al., 2022) A+V 9,288 60 45,867 M 8.6 OmniBench (Li et al., 2024d) A+I 1,142 M 37.8 AV-Odyssey (Gong et al., 2024) A+I 4,555 M 19.5 LongVALE (Geng et al., 2024) A+V 8,400 235 A&M WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2
WorldSense features diverse audio types such as speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench (Gong et al., 2024), OmniBench (Li et al., 2024d)) or feature weak audio-visual correlations (e.g., Video-MME (Fu et al., 2024b)), WorldSense is the first to comprehensively evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a comprehensive benchmark for real-world multimodal understanding assessment.
@@ 4.1 Settings
To comprehensively assess the multi-modal understanding ability, we evaluate three types of MLLMs: (1) open-source audio-visual models, such as Unified-IO-2 (Lu et al., 2024b), OneLLM (Han et al., 2024), and VideoLLaMA2 (Cheng et al., 2024); (ii) open-source MLLMs, such as Qwen2-VL (Wang et al., 2024a), LLaVA-OneVision (Li et al., 2024a), InternVL2.5 (Chen et al., 2024b), LLaVA-Video (Zhang et al., 2024d), and so on; (iii) proprietary MLLMs, such as Claude 3.5 Sonnet (Anthropic, 2024), GPT 4o (Hurst et al., 2024), Gemini 1.5 Pro (Team et al., 2024a), and Gemini 2.5 Pro (Comanici et al., 2025). For all evaluations, we strictly adhere to each model’s official implementation guidelines and the recommended pre-processing procedures. Video frame extraction follows the official configurations specified by corresponding MLLMs, while proprietary models are evaluated according to their API specifications and recommended input formats. Model performance is assessed through direct comparison between model outputs and ground-truth.
@@
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L (Lu et al., 2024b) 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL (Lu et al., 2024b) 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL (Lu et al., 2024b) 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM (Han et al., 2024) 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 Cheng et al. (2024) 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 VITA-1.5 Fu et al. (2025) 7B 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9 Qwen2.5-Omni Xu et al. (2025a) 7B 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4 video-SALMONN 2+ (Tang et al., 2025) 7B 57.1 54.4 48.9 50.9 49.1 51.1 44.9 51.0 50.9 Qwen3-Omni (Xu et al., 2025b) 7B 58.7 60.5 54.5 53.8 55.4 46.8 48.8 52.2 54.0 video-SALMONN 2+ (Tang et al., 2025) 72B 59.0 63.1 54.0 59.9 58.1 54.1 51.9 54.4 56.5 Open-Source Video MLLMs Video-LLaVA (Lin et al., 2023) 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 (Grattafiori et al., 2024) 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL (Wang et al., 2024a) 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 (Ye et al., 2024) 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision (Li et al., 2024a) 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 (Chen et al., 2024b) 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video (Zhang et al., 2024d) 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet (Anthropic, 2024) - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o (Hurst et al., 2024) - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro (Team et al., 2024a) - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0 Gemini 2.5 Flash (Comanici et al., 2025) - 51.8 50.2 54.1 51.2 59.6 50.6 51.6 51.5 52.3 Gemini 2.5 Pro (Comanici et al., 2025) - 64.9 66.0 65.8 68.1 69.7 65.7 63.5 61.3 65.1
@@ 4.2 Results on WorldSense
First, current open-source video models are limited in their performance as they process only visual information. This restriction highlights a significant gap in their ability to perform complex, multi-modal understanding tasks, as evidenced by their maximum performance score of only $54.0\%$. The results underscore the inadequacies of relying solely on visual processing, emphasizing the need to integrate audio inputs for a more comprehensive understanding in practical applications.
@@
Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet demonstrate performance comparable to the leading open-source video MLLMs. Gemini 2.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of $65.1\%$. However, this performance still falls considerably short of requirements for reliable real-world applications, indicating substantial room for improvement.
@@
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Unified-IO-2 L (Lu et al., 2024b) Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL (Lu et al., 2024b) Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL (Lu et al., 2024b) Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM (Han et al., 2024) Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 Cheng et al. (2024) Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 VITA-1.5 Fu et al. (2025) Audio 30.2 35.6 36.3 30.9 32.2 32.2 31.4 33.3 32.9 + Caption 39.2 39.8 37.2 37.5 37.5 35.2 34.9 38.4 37.5+4.6 + Video 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9+4.0 Qwen2.5-Omni Xu et al. (2025a) Audio 40.0 38.2 36.0 33.5 31.1 30.5 32.3 33.3 34.9 + Caption 40.0 37.9 38.9 33.5 36.7 37.8 37.7 38.9 37.9+3.0 + Video 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4+10.5 Gemini 1.5 Pro (Team et al., 2024a) Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4